OpenStudy (anonymous):

In two years, Linda will be half as old as her mother. her mother is 38 years old now. what is lindas current age?

OpenStudy (anonymous):

\[(38\div2)-2\]

OpenStudy (campbell_st):

let linda's current age be x in 2 years time linda will be x + 2 her mother will be 38 + 2 = 40 and linda will be half her mothers' age \[x + 2 = \frac{40}{2}\] now just solve for x ( linda's current age)